a satellite of mass 200 kg is displaced by a distance of 2 Rs gift from a point on the earth surface where R is the radius of the earth calculate the work done on the satellite by the force of gravity take the values of constant is in given your call calculation from instead of physical data given in the course

1
Expert's answer
2022-02-14T12:58:41-0500

"A=-\int_R^{2R}FdR=-\int_R^{2R}\frac{GmMdR}{R^2}={GmM}(\frac 1R-\frac 1{2R})=\frac{GmM}{2R}."